OpenStudy (anonymous):

derivative of f(x)= 56x/(x^2;4)^2

OpenStudy (anonymous):

Use the quotient rule.

OpenStudy (anonymous):

\[ \left(\frac{f}{g}\right)' = \frac{f'g-fg'}{g^2} \]

OpenStudy (anonymous):

yea i did and i keeping getting wrong anser

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

what is \(f\) and \(g\)?

OpenStudy (anonymous):

56x/(x^2+4)^2

OpenStudy (abb0t):

You could also use product rule if you feel more comfortable using that. Remember: \(\frac{ 1 }{ (a+b)^n } = (a+b)^{-n}\)

OpenStudy (anonymous):

I know, so what is \(f'\) and \(g'\)?

OpenStudy (anonymous):

56 and 2x

OpenStudy (anonymous):

Nope, your \(g'\) is wrong.

OpenStudy (anonymous):

You either have to expand it, or use the chain rule.

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

3x^3+32x

OpenStudy (anonymous):

I need to see your steps

OpenStudy (anonymous):

(x^2+4)^2 (x^2+4)(x^2+4) x^4+4x^2+4x^2+16 x^4+8x^2+16 3x^2+16x

OpenStudy (anonymous):

you messed up on \((x^4)'\)

OpenStudy (anonymous):

4x^3

OpenStudy (anonymous):

Alright, now you can use the quotient rule.

OpenStudy (anonymous):

(56)(x^2+4)^2-(56)(4x^3+16)/((x^2+4)^2)^2

OpenStudy (anonymous):

you want 56x on the second term

OpenStudy (anonymous):

ops

OpenStudy (anonymous):

then simplify a bit and you're done.

OpenStudy (anonymous):

simplyging is the problem !